# Convergent Privacy Framework for Multi-layer GNNs through Contractive Message Passing

### v2

@@ Acknowledgment
-A Review on Message Passing GNNs
Every layer in the Graph Convolutional Network (GCN) [iclr/kipf2017semi] can be expressed as a message passing layer, where the aggregation function simply computes weighted sums of the features of the neighbors. The GCN layer applied to graph $\bm{G}$} with a node feature matrix ${\bm{X}}^{(k)}$ can be expressed as:
$\text{GCN}_{G}({\bm{X}}_{u}^{(k)})=\sigma\left(\hat{{\bm{A}}}{\bm{X}}^{(k)}{\bm{W}}^{(k)}\right),$ (9)
where $\hat{{\bm{A}}}={\bm{D}}^{-\frac{1}{2}}({\bm{A}}+{\bm{I}}){\bm{D}}^{-\frac{1}{2}}$ is the symmetrically normalized adjacency matrix of the graph, $\bm{I}$} is the identity matrix, $\bm{D}$} is the degree matrix of the graph (a diagonal matrix where $d_{ii}=\sum_{j}a_{ij}$), $a_{ij}$ is the $(i,j)$-th entry of the adjacency matrix $\bm{A}$}, ${\bm{W}}^{(k)}$ is a learnable weight matrix at layer $k$, and $\sigma$ is a non-linear activation function.
Other simple variants include replacing the mean-type aggregation $\hat{{\bm{A}}}{\bm{X}}^{(k)}$ in GCN with a random walk adjacency matrix $\tilde{{\bm{A}}}{\bm{X}}^{(k)}$, where $\tilde{{\bm{A}}}={\bm{D}}^{-1}{\bm{A}}$ is the random walk normalized adjacency matrix. Alternative aggregation functions include sum or max aggregation, where $\hat{{\bm{A}}}{\bm{X}}^{(k)}$ is replaced with $\sum_{v\in{\mathcal{N}}(u)}{\bm{X}}_{v}^{(k)}$ (like in GAP) or $\max_{v\in{\mathcal{N}}(u)}{\bm{X}}_{v}^{(k)}$, respectively.
@@ Appendix A Preliminary on $f$-DP and Related Definitions
The $f$-differential privacy framework [journal/dong2022gaussian, tit/KairouzOV17] is based on hypothesis testing where $f$ denotes a trade-off function between type I and type II errors. Given two output distributions $P$ and $Q$ of a mechanism $\mathcal{M}$ on neighboring datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$, the problem is to distinguish between them: $H_{0}:\text{dataset is }\mathcal{D}$ vs. $H_{1}:\text{dataset is }\mathcal{D}^{\prime}$. For a rejection function $\phi:\mathcal{X}\to[0,1]$, the type I error is $\mathbb{E}_{P}\phi$ and the type II error is $1-\mathbb{E}_{Q}\phi$. The privacy guarantee is formalized through the trade-off function defined below.
@@
Definition 9 (Trade-off Function [icml/BokShifted24, Theorem 4.2]).
@@
A special case of $f$-DP is the Gaussian differential privacy (GDP) [journal/dong2022gaussian], which is defined as follows.
Definition 10 (Gaussian Differential Privacy [icml/BokShifted24, Definition 2.1]).
@@
Definition 11 ([journal/dong2022gaussian, Definition 3.1]).
@@
The tensor product of trade-off functions satisfies the following properties as proved in [journal/dong2022gaussian, Proposition D.1].
Lemma 2 ([journal/dong2022gaussian, Proposition D.1]).
@@
Theorem 7 (GDP to RDP [icml/BokShifted24, Lemma A.4]).
@@
Theorem 8 (RDP to DP [csf/mironov2017renyi, Proposition 3]).
@@ A-A Preliminaries: Perturbed Message Passing
Recall that each layer of a perturbed message-passing GNN updates the node features following Equation 2, where ${\bm{Z}}\sim\mathcal{N}\bigl(0,\sigma^{2}\mathrm{I}_{d}\bigr)$ is Gaussian noise added to obscure the true output of $\mathrm{MP}_{{\bm{G}}},$ and $\Pi_{\mathcal{K}}$ projects onto some convex feasible set $\mathcal{K}$}. As shown in Proposition 2, for a single layer, the privacy cost depends only on $\Delta(\mathrm{MP})$ and $\sigma.$ However, by the composition theorem of RDP [csf/mironov2017renyi], stacking $K$ such layers yields a cost that scales linearly in $K$ (Corollary 2).
@@
This linear growth of the privacy budget is common in multi-hop settings, such as [uss/sajadmanesh2023gap], but is troublesome for deep GNNs, especially when many message-passing steps are needed for long-range dependencies. Excessively large privacy loss often undermines the model’s utility.
@@
By the Gaussian mechanism guarantee under Rényi differential privacy (RDP) [csf/mironov2017renyi, Corollary 3], the map $X\mapsto\mathrm{MP}_{{\bm{G}}}(X)+Z$ is $\bigl(\alpha,\alpha\,\Delta^{2}(\mathrm{MP})/(2\sigma^{2})\bigr)\text{-RDP}.$ The post-processing theorem then ensures that projecting to $\mathcal{K}$} does not increase the privacy cost. Converting from RDP to $(\epsilon,\delta)\text{-DP}$ [csf/mironov2017renyi, Proposition 3] completes the proof. ∎
@@ A-B Proof of results in Section V-A
The proof is reproduced from the original proof in [icml/BokShifted24], we include it here for self-containedness and to clarify that the contractive condition (Lipschitz constant bounded by $\gamma<1$) is only required for iteration steps $k\geq 2$.
@@
The key idea in the proof, following [icml/BokShifted24], is to construct an auxiliary interpolating sequence $\{\widetilde{{\bm{X}}}^{(k)}\}_{k=0}^{K}$ between the two CNI processes. For each step $k$, we define:
@@
We now recall a lemma from [icml/BokShifted24] that establishes the trade-off function bound for one step of interpolation.
Lemma 4 ([icml/BokShifted24, Lemma 3.2]).
@@
where the second inequality follows from the fact that $\Pi_{{\mathcal{K}}}$ is a projection onto a convex set and hence has Lipschitz constant $1$ ([icml/BokShifted24, Lemma 2.9]). This concludes the proof of the claim by induction. We let $a_{k+1}=\lambda_{k+1}(c_{k+1}z_{k}+s)$ for $k=0,\dots,K-1$. In particular, $a_{1}=\lambda_{1}s$ regardless of the value $c_{1}$ since it is multiplied with $z_{0}=0$. This is the main our observation that the Lipchitz constant of the first iteration does not affect the constant $a_{1}$ and consequently the privacy guarantee.
@@
Lastly, by the assumption that $\phi_{k}$ and $\phi^{\prime}_{k}$ are $\gamma$-Lipschitz for $k=2,\dots,K$, we can let $c_{k}=\gamma$ for $k=2,\dots,K$. We set $\lambda_{k}=\frac{\gamma^{K-k}(1-\gamma^{2})}{1-\gamma^{K-k+2}-\gamma^{k}+\gamma^{K}}$ for $k=1,\dots,K$ as computed in [icml/BokShifted24, Lemma C.6]. Then $z_{k}=\frac{(1-\gamma^{k})(1-\gamma^{K-k})}{(1+\gamma^{K})(1-\gamma)}s$, and consequently, $a_{k}=\frac{\gamma^{K-k}(1+\gamma)}{1+\gamma^{k}}s$. In this case, $\sum_{k=1}^{K}a_{k}^{2}=\frac{(1-\gamma^{K})(1+\gamma)}{(1+\gamma^{K})(1-\gamma)}s^{2}$. This concludes the proof of the theorem. ∎
@@ A-D1 Datasets
Table V lists statistics of these datasets, including Amazon co-purchase networks (Photo, Computers [corr/abs-1811-05868]), social network (Facebook [corr/abs-1102-2166]), and citation networks (Cora, PubMed [sen2008collective]). These datasets are also widely adopted as benchmark datasets to evaluate various GNNs [nips/GuC0SG20, uss/sajadmanesh2023gap, nips/chien2023differentially, nips/YaoJRJ23]. In addition, we use the synthetic graph generation algorithm developed in IGNN [nips/GuC0SG20], to produce chain-structured datasets for convenient and generalized verification of multi-hop aggregations. The chain-structured dataset can be configured with various number of nodes per chain, number of chains per class, and the number of classes. In Table VI, the statistics of the four synthesized chain-structured datasets, namely Chain-S, Chain-M, Chain-L, and Chain-X  are shown. We unify the feature dimension to $5$ and focus on the structure of chain-structured dataset, and we configure the number of nodes to $\{8,10,15\}$ per chain, and the number of chains to $\{3,5\}$ per class.
@@ A-D2 Hardware and software
Our experiments have been conducted on the Ubuntu 20.04.2 LTS server, with AMD Ryzen Threadripper 3970X 32-core CPUs of 256 GB CPU memory and NVIDIA GeForce RTX 3090 of 24GB memory. $\mathsf{CARIBOU}$ equipped with edge- and node-level privacy has been implemented in PyTorch using PyTorch-Geometric (PyG) [corr/abs-1903-02428] framework. DP implementation adopts the autodp library [AutoDP2025], which includes analytical moments accountant and private training with DP-SGD [ccs/AbadiCGMMT016].
@@ A-D5 Model architectures & configurations of $\mathsf{CARIBOU}$
We concatenate $K$ layers ranging from $1$ to $20$, with $K,\alpha_{1},\beta,C_{\mathsf{L}}$ as the hyper-parameters. If the range between two interactive nodes is long, we use larger values to test the influence of hops. Specifically, for Chain-S and Chain-M datasets, $\mathsf{CARIBOU}$ uses the number of hops $K\in\{15,10,9,8,7,6,5,4,3,2,1\}$, while for Chain-S and Chain-M datasets consisting of longer chains, $\mathsf{CARIBOU}$ takes $K\in\{20,17,15,\allowbreak 13,11,9,7,5\}$. As for standard datasets listed in Table V, $K$ is chosen from the union sets of the aforementioned two sets, i.e., $\{20,17,15,\allowbreak 13,11,10,9,8,7,6,5,4,3,2,1\}$. To align with GAP’s configuration [uss/sajadmanesh2023gap], we set the number of hidden units to $\{16,64\}$ and use the SeLU activation function [nips/KlambauerUMH17] at every layer. We adopt the Adam optimizer over $100$ epochs with a learning rate $0.001$, and pick the best accuracy to report.
@@ A-D6 Privacy configuration and parameters
We implement convergent privacy allocation based on GAP’s privacy budget accounting mechanism [uss/sajadmanesh2023gap] and numerically calibrate noise level $\sigma$ by setting $\epsilon$. All edge/node-level private algorithms adopt the Gaussian mechanism to sample the i.i.d. noise to realize the desired $(\epsilon,\delta)$-DP. For comprehensive evaluation, we consider different choices of $\epsilon\in\{1,2,4,8,16,32\}$, while $\delta$ is set to be smaller than the inverse number of edges for EDP or the inverse number of nodes for NDP.
@@ A-G Larger Datasets Evaluation
To further assess the scalability on large real-world graphs, we additionally evaluate $\mathsf{CARIBOU}$ on the Reddit2 dataset [zeng2019graphsaint]. Reddit2 contains over 232K nodes and 23M edges, representing a significantly larger and structurally richer benchmark. Table XIV summarizes the top-1 accuracy under both EDP and NDP across privacy budgets $\epsilon\in\{1,2,4,8,16,32\}$.
@@ A-I1 Differential privacy for graph structures
Euclidean data clearly states which data points are associated with a particular individual, for example, tabular data. Unlike tabular data, graph association can be interpreted into the combination by edges and nodes. Corresponding to instance-level DP, the “instance” of graph data can be an edge or a node, naturally called edge DP and node DP. Early stage works [tcc/GehrkeLP11, icdm/MirW09, pakdd/WangWW13, vldb/ChenFYD14, kdd/LuM14] started by private statistics estimation or counting, or release of private graphs under edge DP. Node DP protects against revealing the presence or absence of an individual node along with all its adjacent edges. Prior works consider various node-privacy algorithms [tcc/KasiviswanathanNRS13, pakdd/IftikharW21, sigmod/DayLL16, sigmod/XiaCKHT021, uss/ImolaMC21] tailored to specific graph statistics constrained by bounded node degrees. Node DP typically offers a stronger level of protection compared to edge privacy, as inclusion of a node implies the inclusion of some particular edges. Node DP is also considered harder to achieve than edge DP if maintaining reasonable utility loss.
@@ A-I2 Differentially private GNNs
Sharing model updates can lead to privacy risks, as adversaries may reconstruct training data. Such risks [sp/CarliniCN0TT22, ccs/0001MMBS22] stem from overfitting, where neural networks memorize training data [zhang2021understanding]. Graph structures, which encode relational information, are widely used in applications like intrusion detection [eurosp/ApruzzeseLS23, sp/SommerP10], social recommendation [icde/XiaSH0XP23], and drug discovery [kdd/ZhongM23]. Graph neural networks (GNNs) have emerged as a key approach for learning over graph-structured data, but their message-passing mechanisms can leak sensitive information about nodes and their neighbors. To address these risks, several works [arxiv/sajadmanesh2020differential, ccs/sajadmanesh2021locally, uss/sajadmanesh2023gap, tcc/GuptaRU12] propose differentially private GNNs [arxiv/DaigavaneNode21, corr/abs-2109-08907, uss/sajadmanesh2023gap]. This necessitates injecting calibrated noise after each message passing layer, where the noise scale is proportional to $K$ for a fixed privacy budget $\epsilon$. For instance, Wu et al. [sp/Wulinkteller22] achieve edge-level privacy by adding Laplace noise to adjacency matrix entries under a small DP budget. Kolluri et al. [ccs/KolluriBHS22] improve privacy-utility trade-offs by decoupling graph structure from the neural network architecture, querying the graph only when necessary. However, leveraging contractive hidden states in graph learning for enhanced privacy analysis remains an underexplored avenue.

### v3

@@ Appendix A Preliminary on $f$-DP and Related Definitions
The $f$-differential privacy framework [17, 32] is based on hypothesis testing where $f$ denotes a trade-off function between type I and type II errors. Given two output distributions $P$ and $Q$ of a mechanism $\mathcal{M}$ on neighboring datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$, the problem is to distinguish between them: $H_{0}:\text{dataset is }\mathcal{D}$ vs. $H_{1}:\text{dataset is }\mathcal{D}^{\prime}$. For a rejection function $\phi:\mathcal{X}\to[0,1]$, the type I error is $\mathbb{E}_{P}\phi$ and the type II error is $1-\mathbb{E}_{Q}\phi$. The privacy guarantee is formalized through the trade-off function defined below.
@@
Definition 9 (Trade-off Function [6, Theorem 4.2]).
@@
A special case of $f$-DP is the Gaussian differential privacy (GDP) [17], which is defined as follows.
Definition 10 (Gaussian Differential Privacy [6, Definition 2.1]).
@@
Definition 11 ([17, Definition 3.1]).
@@
The tensor product of trade-off functions satisfies the following properties as proved in [17, Proposition D.1].
Lemma 2 ([17, Proposition D.1]).
@@
Theorem 7 (GDP to RDP [6, Lemma A.4]).
@@
Theorem 8 (RDP to DP [49, Proposition 3]).
@@ A-A Preliminaries: Perturbed Message Passing
Recall that each layer of a perturbed message-passing GNN updates the node features following Equation 2, where ${\bm{Z}}\sim\mathcal{N}\bigl(0,\sigma^{2}\mathrm{I}_{d}\bigr)$ is Gaussian noise added to obscure the true output of $\mathrm{MP}_{{\bm{G}}},$ and $\Pi_{\mathcal{K}}$ projects onto some convex feasible set $\mathcal{K}$}. As shown in Proposition 2, for a single layer, the privacy cost depends only on $\Delta(\mathrm{MP})$ and $\sigma.$ However, by the composition theorem of RDP [49], stacking $K$ such layers yields a cost that scales linearly in $K$ (Corollary 2).
@@
This linear growth of the privacy budget is common in multi-hop settings, such as [59], but is troublesome for deep GNNs, especially when many message-passing steps are needed for long-range dependencies. Excessively large privacy loss often undermines the model’s utility.
@@
By the Gaussian mechanism guarantee under Rényi differential privacy (RDP) [49, Corollary 3], the map $X\mapsto\mathrm{MP}_{{\bm{G}}}(X)+Z$ is $\bigl(\alpha,\alpha\,\Delta^{2}(\mathrm{MP})/(2\sigma^{2})\bigr)\text{-RDP}.$ The post-processing theorem then ensures that projecting to $\mathcal{K}$} does not increase the privacy cost. Converting from RDP to $(\epsilon,\delta)\text{-DP}$ [49, Proposition 3] completes the proof. ∎
@@ A-B Proof of results in Section V-A
The proof is reproduced from the original proof in [6], we include it here for self-containedness and to clarify that the contractive condition (Lipschitz constant bounded by $\gamma<1$) is only required for iteration steps $k\geq 2$.
@@
The key idea in the proof, following [6], is to construct an auxiliary interpolating sequence $\{\widetilde{{\bm{X}}}^{(k)}\}_{k=0}^{K}$ between the two CNI processes. For each step $k$, we define:
@@
We now recall a lemma from [6] that establishes the trade-off function bound for one step of interpolation.
Lemma 4 ([6, Lemma 3.2]).
@@
where the second inequality follows from the fact that $\Pi_{{\mathcal{K}}}$ is a projection onto a convex set and hence has Lipschitz constant $1$ ([6, Lemma 2.9]). This concludes the proof of the claim by induction. We let $a_{k+1}=\lambda_{k+1}(c_{k+1}z_{k}+s)$ for $k=0,\dots,K-1$. In particular, $a_{1}=\lambda_{1}s$ regardless of the value $c_{1}$ since it is multiplied with $z_{0}=0$. This is the main our observation that the Lipchitz constant of the first iteration does not affect the constant $a_{1}$ and consequently the privacy guarantee.
@@
Lastly, by the assumption that $\phi_{k}$ and $\phi^{\prime}_{k}$ are $\gamma$-Lipschitz for $k=2,\dots,K$, we can let $c_{k}=\gamma$ for $k=2,\dots,K$. We set $\lambda_{k}=\frac{\gamma^{K-k}(1-\gamma^{2})}{1-\gamma^{K-k+2}-\gamma^{k}+\gamma^{K}}$ for $k=1,\dots,K$ as computed in [6, Lemma C.6]. Then $z_{k}=\frac{(1-\gamma^{k})(1-\gamma^{K-k})}{(1+\gamma^{K})(1-\gamma)}s$, and consequently, $a_{k}=\frac{\gamma^{K-k}(1+\gamma)}{1+\gamma^{k}}s$. In this case, $\sum_{k=1}^{K}a_{k}^{2}=\frac{(1-\gamma^{K})(1+\gamma)}{(1+\gamma^{K})(1-\gamma)}s^{2}$. This concludes the proof of the theorem. ∎
@@ A-D1 Datasets
Table V lists statistics of these datasets, including Amazon co-purchase networks (Photo, Computers [62]), social network (Facebook [71]), and citation networks (Cora, PubMed [61]). These datasets are also widely adopted as benchmark datasets to evaluate various GNNs [25, 59, 13, 89]. In addition, we use the synthetic graph generation algorithm developed in IGNN [25], to produce chain-structured datasets for convenient and generalized verification of multi-hop aggregations. The chain-structured dataset can be configured with various number of nodes per chain, number of chains per class, and the number of classes. In Table VI, the statistics of the four synthesized chain-structured datasets, namely Chain-S, Chain-M, Chain-L, and Chain-X  are shown. We unify the feature dimension to $5$ and focus on the structure of chain-structured dataset, and we configure the number of nodes to $\{8,10,15\}$ per chain, and the number of chains to $\{3,5\}$ per class.
@@ A-D2 Hardware and software
Our experiments have been conducted on the Ubuntu 20.04.2 LTS server, with AMD Ryzen Threadripper 3970X 32-core CPUs of 256 GB CPU memory and NVIDIA GeForce RTX 3090 of 24GB memory. $\mathsf{CARIBOU}$ equipped with edge- and node-level privacy has been implemented in PyTorch using PyTorch-Geometric (PyG) [23] framework. DP implementation adopts the autodp library [74], which includes analytical moments accountant and private training with DP-SGD [1].
@@ A-D5 Model architectures & configurations of $\mathsf{CARIBOU}$
We concatenate $K$ layers ranging from $1$ to $20$, with $K,\alpha_{1},\beta,C_{\mathsf{L}}$ as the hyper-parameters. If the range between two interactive nodes is long, we use larger values to test the influence of hops. Specifically, for Chain-S and Chain-M datasets, $\mathsf{CARIBOU}$ uses the number of hops $K\in\{15,10,9,8,7,6,5,4,3,2,1\}$, while for Chain-S and Chain-M datasets consisting of longer chains, $\mathsf{CARIBOU}$ takes $K\in\{20,17,15,\allowbreak 13,11,9,7,5\}$. As for standard datasets listed in Table V, $K$ is chosen from the union sets of the aforementioned two sets, i.e., $\{20,17,15,\allowbreak 13,11,10,9,8,7,6,5,4,3,2,1\}$. To align with GAP’s configuration [59], we set the number of hidden units to $\{16,64\}$ and use the SeLU activation function [37] at every layer. We adopt the Adam optimizer over $100$ epochs with a learning rate $0.001$, and pick the best accuracy to report.
@@ A-D6 Privacy configuration and parameters
We implement convergent privacy allocation based on GAP’s privacy budget accounting mechanism [59] and numerically calibrate noise level $\sigma$ by setting $\epsilon$. All edge/node-level private algorithms adopt the Gaussian mechanism to sample the i.i.d. noise to realize the desired $(\epsilon,\delta)$-DP. For comprehensive evaluation, we consider different choices of $\epsilon\in\{1,2,4,8,16,32\}$, while $\delta$ is set to be smaller than the inverse number of edges for EDP or the inverse number of nodes for NDP.
@@ A-G Larger Datasets Evaluation
To further assess the scalability on large real-world graphs, we additionally evaluate $\mathsf{CARIBOU}$ on the Reddit2 dataset [93]. Reddit2 contains over 232K nodes and 23M edges, representing a significantly larger and structurally richer benchmark. Table XIV summarizes the top-1 accuracy under both EDP and NDP across privacy budgets $\epsilon\in\{1,2,4,8,16,32\}$.
@@ A-I1 Differential privacy for graph structures
Euclidean data clearly states which data points are associated with a particular individual, for example, tabular data. Unlike tabular data, graph association can be interpreted into the combination by edges and nodes. Corresponding to instance-level DP, the “instance” of graph data can be an edge or a node, naturally called edge DP and node DP. Early stage works [24, 48, 73, 12, 45] started by private statistics estimation or counting, or release of private graphs under edge DP. Node DP protects against revealing the presence or absence of an individual node along with all its adjacent edges. Prior works consider various node-privacy algorithms [34, 30, 16, 87, 31] tailored to specific graph statistics constrained by bounded node degrees. Node DP typically offers a stronger level of protection compared to edge privacy, as inclusion of a node implies the inclusion of some particular edges. Node DP is also considered harder to achieve than edge DP if maintaining reasonable utility loss.
@@ A-I2 Differentially private GNNs
Sharing model updates can lead to privacy risks, as adversaries may reconstruct training data. Such risks [7, 90] stem from overfitting, where neural networks memorize training data [94]. Graph structures, which encode relational information, are widely used in applications like intrusion detection [5, 63], social recommendation [86], and drug discovery [101]. Graph neural networks (GNNs) have emerged as a key approach for learning over graph-structured data, but their message-passing mechanisms can leak sensitive information about nodes and their neighbors. To address these risks, several works [57, 58, 59, 26] propose differentially private GNNs [15, 52, 59]. This necessitates injecting calibrated noise after each message passing layer, where the noise scale is proportional to $K$ for a fixed privacy budget $\epsilon$. For instance, Wu et al. [78] achieve edge-level privacy by adding Laplace noise to adjacency matrix entries under a small DP budget. Kolluri et al. [38] improve privacy-utility trade-offs by decoupling graph structure from the neural network architecture, querying the graph only when necessary. However, leveraging contractive hidden states in graph learning for enhanced privacy analysis remains an underexplored avenue.